OpenStudy (anonymous):

The minute hand of a clock is 6" long. How far does the tip of the hand move in 35 minutes?

hero (hero):

Hint: \[\frac{35}{60} = \frac{\theta}{360}\] 2nd Hint: \[s = r \theta\] in radians